# The Cone Restriction: An old Approach Revisited

### v5

@@ 1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, Lemma 2.1 and Corollary 2.7 gives
@@
where $\widetilde{f_{S,\mathcal{T}}}:={\left(\widetilde{f_{S}}\right)}_{\mathcal{T}}$
@@
Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. From (2.12),
@@
Fix $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$ and $B\in{\mathcal{B}}_{O_{j}}$. By Lemma 2.22, there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b,B\right)$ for each $b\in{\mathfrak{b}}_{B,O_{j}}$ such that
${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
@@
$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, let
@@
[…]
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.15)
Combine (3.15) and (3.3),
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j+1\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$
The induction closes if $\overline{C}$ and $d$ is large enough. For a ${\rho}^{-1/2}$-sector $\theta$ with $1\leq\rho\leq{\rho}_{j+1}$, Lemma 2.6, Lemma 2.27, and Corollary 2.2 gives
@@
${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}$ (3.16)
Combine (3.16) and (3.4),
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
[…]
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.17)
Combine (3.5) and (3.17)
@@
Since $\overline{C}$ and $d$ is large enough, one has
@@
Like the cellular case, ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$, and thus the induction closes.
Remark 3.1.
A trivial observation is
$\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
Hence, the accumulation of error is still negligible after the algorithm terminates. We omit the details.
Let $D:=d^{c\left(J\right)}$. If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.19)
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.20)
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.21)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.22)
${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.23)
@@
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.24)
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.25)
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.26)
3.2 The second algorithm
[alg 2] is established by repeated application of [alg 1] and terminates if [tiny] term dominates. Consider a family of Lebesgue exponents $p_{i}$ for $k\leq l\leq n$, which will be determined in Section 4, satisfying
$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$
and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ in term of the $p_{i}$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
@@ 3.2 The second algorithm
Here $A\in\mathbb{N}$ is a large constant.
Output: The algorithm outputs the following objects.
$\bullet$ A collection $\mathcal{O}$ of cells such that each $O\in\mathcal{O}$ is included in a $R^{O\left({\delta}_{0}\right)}$-ball.
$\bullet$ For each $O\in\mathcal{O}$ a function $f_{O}$.
$\bullet$ Scales $\overrightarrow{r}=\left(r_{n},\dots,r_{m-1}\right)$ satisfying $R=r_{n}>r_{n-1}>\dots>r_{m-1}=1$.
$\bullet$ Integers $\overrightarrow{A}=\left(A_{n},\dots,A_{m-1}\right)$ satisfying $A=A_{n}>A_{n-1}>\dots>A_{m-1}$.
$\bullet$ A tuple of parameters $\overrightarrow{D}=\left(D_{n},\dots,D_{m-1}\right)$.
$\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain. Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has a multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and a bounded degree. Also, for $l\leq s\leq t\leq n$:
$Z_{s}=Z^{\prime}_{s}+b\mathrm{,\ \ for\ some\ }b\mathrm{\ with\ }\left|b\right|\lesssim r^{1/2+{\delta}_{t}}_{t}\ \ \mathrm{and\ }Z^{\prime}_{s}\subseteq Z_{t}$
$\bullet$ For $m\leq l\leq n$ an assignment of a function $f_{{\overrightarrow{S}}_{l}}$ to each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$. Each $f_{{\overrightarrow{S}}_{l}}$ is tangent to $S_{l}=\left(Z_{l},B_{r_{l}}\right)$, the first component of ${\overrightarrow{S}}_{l}$.
@@
2.There holds
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D^{1+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.31)
@@
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.32)
@@
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.33)
@@
$\left(n+2-l\right)$-th step. Suppose the algorithm has run through $\left(n+1-l\right)$ steps and the following things have been determined.
@@
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$. Let $\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and $\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote $\mathrm{a}\left(J\right)$ and $\mathrm{c}\left(J\right)$ output by [alg 1] respectively. Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively. Let
@@
[…]
@@
By (3.18) and (3.34), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
@@
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$, let $\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ denote the final cells output by [alg 1]. For each $O\in\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ let $f_{O}$ denote the function corresponding to $O$ output by [alg 1]. Finally, [alg 2] outputs the following additional things, and then terminates.
$m=l,\ \ r_{m-1}:=1,\ \ \mathcal{O}:=\bigcup_{{\overrightarrow{S}}_{m}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{m\mathrm{,tiny}}}{\mathcal{O}\left[{\overrightarrow{S}}_{m}\right]}$
[…]
Tangent-dominated case If (3.34) fails, then
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.36)
By (3.18) and (3.36), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
@@
$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$, let $S\left[{\overrightarrow{S}}_{l}\right]$ denote the grains output by [alg 1]. For each $S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]$, let ${\overrightarrow{S}}_{l-1}:=\left(S_{l\mathrm{-}\mathrm{1}},{\overrightarrow{S}}_{l}\right)$ and $f_{{\overrightarrow{S}}_{l-1}}$ denote the function corresponding to $S_{l\mathrm{-}\mathrm{1}}$ output by [alg 1]. Finally, we define
@@
Remark 3.2.
If $m<k$, then Lemma 2.21 implies
which contradicts (3.28). Thus, $m\geq k$.
[…]

### v7

@@ 3.1. Algorithm 1
Moreover, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$ and $S=(\mathbf{Y},B)\in\mathcal{S}$ with $B\in\mathcal{B}_{O_{j}}$, Lemma 2.1 together with Corollary 2.7 yields
@@
where $\big(f_{S,\mathcal{T}}\big)^{\widetilde{}}:={\left(f_{S}\right)}_{\mathcal{T}}^{\widetilde{}}$
@@
Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. By (2.5),
@@
Fix $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$ and $B\in{\mathcal{B}}_{O_{j}}$. Arguing as in [6], for each $b\in{\mathfrak{b}}_{B,O_{j}}$ there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\bm{\mathrm{Z}}}_{O_{j}}+b,B\right)$ such that
${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
@@
$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, set
@@
We now verify the properties. Properties 1 and 2 follow from [11]; here we verify Properties 3 and 4.
Property 3. By Lemma 2.18,
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$
Combining this with (3.3), we obtain
$\|f_{O_{j+1}}\|_{2}^{2}\lesssim r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j+1)(m-1)}\|f\|_{2}^{2}.$
The induction closes provided $\overline{C}$ and $d$ are chosen sufficiently large. For a $\rho^{-1/2}$-sector $\theta$ with $1\leq\rho\leq\rho_{j+1}$, Lemma 2.6, Lemma 2.19, and Corollary 2.2 yield
@@
$\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim r^{O(\delta_{m})}\left(\frac{\rho_{j+1}}{\rho_{j}}\right)^{-(n-m)/2}\|f_{O_{j}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta^{\prime})}$
Combining this and (3.4), we obtain
$\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim\\ r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}$
Since $\overline{C}$ and $d$ are chosen sufficiently large, this verifies Property 3.
Property 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$, Lemma 2.19 implies
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$
Combining this with (3.5) we have
@@
Since $\overline{C}$ and $d$ are chosen sufficiently large, one has
@@
As in the cellular case, we have $\uparrow^{j}(\uparrow\mathcal{T})\subseteq\uparrow^{j+1}\mathcal{T}$, and thus the induction closes.
Summary. We now summarize the output in a concise form. Since
$\mathrm{c}\left(J\right)\lesssim\log{r},\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}\log{{\delta}^{-1}_{m-1}}$ (3.15)
the accumulated error remains negligible after the algorithm terminates. Let $D:=d^{c\left(J\right)}$. If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.16)
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.17)
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.18)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.19)
${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.20)
@@
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.21)
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.22)
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.23)
3.2. Algorithm 2
Algorithm 2 is obtained by repeatedly applying Algorithm 1 and terminates when the [tiny] case dominates. Consider a family of Lebesgue exponents $p_{i}$ for $k\leq l\leq n$, which will be determined in Section 4, satisfying
$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$ (3.24)
and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
@@ 3.2. Algorithm 2
Here, $A\in\mathbb{N}$ is chosen to be sufficiently large.
Output. The algorithm produces the following objects:
  • A collection $\mathcal{O}$ of cells, each contained in a $R^{O(\delta_{0})}$-ball.
  • For each $O\in\mathcal{O}$, a function $f_{O}$.
  • Scales $\overrightarrow{r}=(r_{n},\dots,r_{m-1})$ satisfying $R=r_{n}>r_{n-1}>\cdots>r_{m-1}=1$.
  • Integers $\overrightarrow{A}=(A_{n},\dots,A_{m-1})$ satisfying $A=A_{n}>A_{n-1}>\cdots>A_{m-1}$.
  • A tuple of parameters $\overrightarrow{D}=(D_{n},\dots,D_{m-1})$.
  • For $m\leq l\leq n$, a family $\overrightarrow{\mathcal{S}}_{l}$ of level (nl+1)(n-l+1) multigrains. Each $\overrightarrow{S}_{l}=((Z_{l},B_{r_{l}}),\dots,(Z_{n},B_{r_{n}}))\in\overrightarrow{\mathcal{S}}_{l}$ has multiscale $\overrightarrow{r}_{l}=(r_{l},\dots,r_{n})$ and bounded degree. Moreover, for $l\leq s\leq t\leq n$,
    $Z_{s}=Z^{\prime}_{s}+b\quad\text{for some }b\text{ with }|b|\lesssim r_{t}^{1/2+\delta_{t}},\qquad Z^{\prime}_{s}\subseteq Z_{t}.$
  • For $m\leq l\leq n$, an assignment of a function $f_{\overrightarrow{S}_{l}}$ to each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$. Each $f_{\overrightarrow{S}_{l}}$ is tangent to $S_{l}=(Z_{l},B_{r_{l}})$, the first component of $\overrightarrow{S}_{l}$.
@@
2.The inequality
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.29)
holds.
@@
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.30)
@@
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.31)
@@
$\left(n+2-l\right)$-th step. Suppose the algorithm has run for (n+1l)(n+1-l) steps and the following things have been determined.
@@
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$, we apply Algorithm 1 to the $k$-broad norm $\|Ef_{\overrightarrow{S}_{l}}\|_{BL^{p_{l}}_{k,A_{l}}(B_{r_{l}})}$. Let $r[\overrightarrow{S}_{l}]$, $D[\overrightarrow{S}_{l}]$, and $A[\overrightarrow{S}_{l}]$ denote the quantities $r_{J}$, $D$, and $A_{J}$ returned by Algorithm 1, respectively. Let
@@
Consequently, at least one of the following two cases must hold:
Tiny-dominated Case. We say that the tiny-dominated case happens if
@@
By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ and $D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
@@
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tiny}}$, let $\mathcal{O}[\overrightarrow{S}_{l}]$ denote the final cells produced by Algorithm 1. For each $O\in\mathcal{O}[\overrightarrow{S}_{l}]$, let $f_{O}$ be the corresponding function returned by Algorithm 1. Algorithm 2 then outputs the following additional data and terminates:
$m=l,\qquad r_{m-1}:=1,\qquad\mathcal{O}:=\bigcup_{\overrightarrow{S}_{m}\in\overrightarrow{\mathcal{S}}^{\prime}_{m,\mathrm{tiny}}}\mathcal{O}[\overrightarrow{S}_{m}].$
Tangent-dominated Case. If (3.32) fails, then
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$
By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
@@
$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tang}}$, let $S[\overrightarrow{S}_{l}]$ denote the collection of grains produced by Algorithm 1. For each $S_{l-1}\in S[\overrightarrow{S}_{l}]$, define
$\overrightarrow{S}_{l-1}:=(S_{l-1},\overrightarrow{S}_{l}),$
and let $f_{\overrightarrow{S}_{l-1}}$ denote the corresponding function returned by Algorithm 1. Finally, we define